# Ultrastable jammed sphere packings with a wide range of particle dispersities

### v1

@@ III Results
In systems prepared without SWAP and TDOF moves, $Z_{\rm J}=5.38$ for $\Delta=0$, and decreases monotonically with increasing $\Delta$, to a minimum of $Z_{\rm J}=2.19$ for $\Delta=0.50$. Over the same range of $\Delta$, the rattler fraction increases monotonically from 7.4% to 67%. Both of these trends arise because (as discussed in Section II) our LS-like growth algorithm terminates and systems are considered jammed when $\tilde{\mathcal{G}}$ drops to zero within numerical precision, i.e. when the particles within the percolating, mechanically-rigid backbones whose formation normally coincides with jamming Moukarzel (1998) can no longer grow without producing overlaps. As $\Delta$ increases, these backbones are increasingly dominated by the large particles Monti et al. (2022, 2023), and thus the jammed states contain an increasing number of small rattlers. We emphasize, however, that these results are not artifacts of employing a LS-like particle-growth algorithm; the same trends are present in jammed Hookean-sphere packings prepared via compression at constant pressure Monti et al. (2022, 2023).
[…]
Relatively few studies of jamming have examined hyperstatic packings, i.e. systems with $Z_{\rm J}>Z_{\rm iso}$. Brito et al.  found that SWAP dynamics increase the number of contacts required to mechanically stabilize polydisperse packings, with $\delta Z=Z_{\rm J}-Z_{\rm iso}\sim\Delta^{1/2}$ when rattlers are removed before calculating $Z_{\rm J}$ Brito et al. (2018). It is reasonable to suppose that further increases in $Z_{\rm J}$ are made possible by the TDOF moves, and we will argue below that this is indeed the case. However, we emphasize that the connections between the hyperstaticity and mechanical stability of high-$\Delta$ packings are far less clear than is the case for their low-$\Delta$ counterparts. For example, we find that the fraction of particles with $Z_{i}>12$ (which is impossible for monodisperse hard spheres) is nonzero for all $\Delta>0.12$ ($\Delta>0.11$), and exceeds 1% for all $\Delta>0.37$ ($\Delta>0.21$) in systems prepared without (with) SWAP and TDOF moves. The emergence of long tails in the $P(Z_{i})$ distributions for $\Delta>0.20$ indicates that caution must be exercised when using $Z_{\rm J}$ as a metric to assess high-$\Delta$ packings’ mechanical stability.
@@
Further examination of the $P(Z_{i})$ distributions clarifies the origin of $Z_{\rm J}(\Delta)$’s very-strong protocol-dependence. Systems prepared without SWAP and TDOF moves have many “floaters” (particles with $Z_{i}=0$), and also substantial populations of rattlers with contacts (particles with $1\leq Z_{i}\leq 3$). Their $P(Z_{i})$ increase rapidly with increasing $\Delta$ for all $0\leq Z_{i}\leq 2$, and slowly with increasing $\Delta$ for $Z_{i}=3$. For all $Z_{i}>4$, their $P(Z_{i})$ decrease monotonically with increasing $\Delta$, consistent with the decreasing $Z_{\rm J}(\Delta)$ shown in panel (a). Minimally-mechanically-stable particles with $Z_{i}=4\equiv d+1$ (where $d$ is the spatial dimension non ) exhibit a qualitatively-different trend; $P(4)$ is nonomonotonic in $\Delta$, peaking at $\Delta\simeq 0.27$.
@@
The shape of the $P(Z_{i})$ distributions also changes qualitatively as polydispersity increases. For all $\Delta\leq 0.13$, $P(5)>P(6)>P(4)$. Then, for $0.14\leq\Delta\leq 0.20$, $P(5)>P(4)>P(6)$. Finally, for all $\Delta>0.20$, $P(4)>P(5)>P(6)$. This crossover from dominance of locally-isostatic particles with $Z_{i}=6$ to minimally-mechanically-stable particles with $Z_{i}=4$ has not (to the best of our knowledge) been previously reported. The emergence of the peak at $Z_{i}=4$ coincides with the emergence of long high-$Z_{i}$ tails in the $P(Z_{i})$ that correspond to large spheres contacted by many smaller spheres; related issues were discussed in Refs. Monti et al. (2022, 2023).
@@
In systems prepared with SWAP and TDOF moves, the $P(Z_{i})$ are radically different. First, all floaters and rattlers with $Z_{i}\leq 2$ are eliminated. The remaining rattlers all have $Z_{i}=3$, and the $P(3)$ are 2-5 times lower than in systems prepared without these moves for all $\Delta$. The $P(4)$ are lower for all $\Delta<0.15$, but higher for all $\Delta\geq 0.15$. Finally, for all $Z_{i}>4$, the $P(Z_{i})$ are larger and decrease slower with increasing $Z_{i}$, particularly for $\Delta\gtrsim 0.2$. All of these differences can be explained as follows: LS-like particle-growth algorithms, with or without SWAP moves, terminate with large numbers of small floaters/rattlers occupying the gaps between the large particles forming the jammed backbone. Adding TDOF moves allows these particles to continue growing until most of them also jam. We believe that the remaining rattlers with $Z_{i}=3$ indicate that the algorithm described in Sec. II is imperfect; improving it might produce ideally-stable, rattler-free packings bol .
@@
Figure 4: Typical jammed states for $N=1000$. The top and bottom rows respectively show packings generated without and with SWAP and TDOF moves. Columns from left to right show Δ=0.10, 0.20, 0.30, 0.40,and $\Delta=0.10,\ 0.20,\ 0.30,\ 0.40,\ \rm{and}\ 0.50$. Particle colors range from purple to red in order of increasing $\sigma$. Only spheres whose centers lie within the boxes are shown.
@@
One might expect that the large changes in $Z_{\rm J}(\Delta)$ and $P(Z_{i})$ enabled by employing TDOF moves would necessarily coincide with large changes in the jammed states’ $P(\sigma)$. The latter would cast doubt on the utility of our results since TDOF moves are unphysical. We find, however, that no such changes occur. As illustrated in Figure 3, the changes in $P(\sigma)$ associated with employing TDOF moves are negligible for $0<\Delta\lesssim 0.45$. For $\Delta=0$, when the moves are not employed, $P(\sigma)=\delta(\sigma-1)$. When they are, $P(\sigma)$ is nonzero over a small range .$1.04.99<\sigma<$ 1.04; this is only a very small broadening. For $\Delta\gtrsim 0.45$, $P(\sigma)$ decreases slightly for $\sigma\simeq\sigma_{\rm min}$ because the smallest particles are most able to grow as they fill in the jammed backbone’s gaps. We believe that this effect is responsible for the increase of $\phi_{\rm J}(\Delta)$ above its linear-fit value (Fig. 1) and the hyperstatic $Z_{\rm J}(\Delta)>Z_{\rm iso}$ (Fig. 2a).
@@
The TDOF moves employed in Refs. Hagh et al. (2022); bol ; Kapteijns et al. (2019) produced much larger changes in $P(\sigma)$ because they were applied throughout the particle-growth process. In contrast, our approach minimizes changes in $P(\sigma)$ by not implementing TDOF moves until near the end of this process. Thus one can imagine that the effects of including TDOF moves highlighted in Fig. 2 could be at least partially reproduced in experiments using a suitably designed packing-preparation protocol.
Previous studies of how preparation protocol affects jammed packings’ ($\phi_{\rm J},\Psi$) Donev et al. (2004); Ozawa et al. (2017) have shown that for fixed $P(\sigma)$ and $\Delta$, systems with larger $\phi_{\rm J}$ have higher rattler fractions $f_{\rm ratt}$, more bond-orientational and icosahedral-like order, reduced Voronoi-cell asphericity, and larger low-$k$ spectral densities (i.e. smaller long-wavelength packing-fraction fluctuations Zachary et al. (2011)). We have already shown that employing TDOF moves reverses the $\phi_{\rm J}$-dependence of $f_{\rm ratt}$, and we do not attempt to duplicate the other structural analyses here. Instead we ask a simpler question: does employing SWAP and TDOF moves during particle growth qualitatively change packings’ structural order, in ways that are readily apparent to the naked eye? Figure 4 shows that the changes are, in fact, rather subtle. In particular, employing SWAP moves does not produce fractionated polycrystalline packings, as it did for low-$\alpha$ ellipses Hoy (2024) and for hard-sphere liquids with $P(\sigma)\sim\sigma^{-3}$ and $\Delta\lesssim 0.10$ that were equilibrated at sufficiently large $\phi_{\rm eq}$ Ninarello et al. (2017). The only difference that is clear in these snapshots is that employing SWAP and TDOF moves allows small particles to more efficiently fill the gaps between the larger ones.
@@
Figure 5: Protocol dependence of pair correlation functions for $N=1000$. Solid and dashed curves show results obtained with and without SWAP, and the inset highlights the “correlation hole” effect for $\Delta=0.10$ and $0.20$. All lengths have been scaled to yield $\langle\sigma\rangle=1$.
@@
Examining the packings’ pair correlation functions $g(r)$ yields additional insights. As illustrated in Figure 5, employing SWAP and TDOF moves can produce packings with better-organized coordination shells. In particular, it yields a clearer separation between nearest and second-nearest neighbors, as indicated by the deeper first minima and higher second maxima of $g(r)$. These differences, however, only occur for $\Delta\lesssim 0.2$. For $\Delta\gtrsim 0.3$, the $g(r)$ are nearly protocol-independent. The increasingly-wide $P(\sigma)$ distributions (Fig. 3) make the first and second coordination shells indistinguishable, and the $g(r)$ are nearly ideal-gas-like for $r>\sigma_{\rm max}$.
@@
Finally, to demonstrate that the various structural differences discussed above strongly affect the packings’ mechanical properties, we examine their bulk moduli $K(\Delta)$. Multiple previous studies have examined how $K$ depends on the pair potential, particle size ratio (in bidisperse systems), and the overcompression $\Delta\phi=\phi-\phi_{\rm J}$ O’Hern et al. (2003); Mizuno et al. (2016); Goodrich et al. (2016); Petit et al. (2022). Here we focus on how $K$ depends on $\Delta$ and preparation protocol. For Hertzian interactions, $P\sim(\Delta\phi)^{3/2}$ in the small-$\Delta\phi$, elastic regime where systems deform affinely O’Hern et al. (2003). The resulting bulk moduli are singular at jamming, i.e. $K\sim(\Delta\phi)^{1/2}$, but a useful measure of their relative magnitudes can be obtained by writing
@@
Figure 6 shows the $K_{0}$ values obtained by fitting the $P(\Delta\phi)$ data to $P=P_{0}(\Delta\phi)^{3/2}$ over the power-law scaling regime, i.e. over the range $10^{-4}\leq\Delta\phi\leq 10^{-2}$, and then plugging the obtained $P_{0}$ values and results for $\phi_{\rm J}$ discussed above into Eq. 7.
Figure 6: Protocol dependence of $K_{0}$ values for $N=1000$. All lengths have been scaled to yield $\langle\sigma\rangle=1$, and $K_{0}$ values are given in urits of $\varepsilon/\langle\sigma\rangle^{3}$. The dashed line shows a linear fit to $K_{0}$ over the range $0.25\leq\Delta\leq 0.45$, and the inset shows the fractional increase in $K_{0}$ obtained by employing SWAP and TDOF moves.
@@
is the classical Born term arising from affine compression and $K_{\rm na}$ captures the reduction of $K$ by nonaffine particle displacements Maloney and Lemaître (2004); Mizuno et al. (2016). The $1/V$ prefactor and the purely-repulsive interactions [i.e., $U(r_{ij})=0$ for $r_{ij}>\sigma_{ij}$] respectively imply $K\sim\phi_{\rm J}$ and $K\sim Z_{\rm J}$ for monodisperse systems. Polydisperse systems are more complicated, as the $Z_{\rm J}$-dependence of their bulk moduli depends sensitively on which particles are in contact.
Our results suggest that as $\Delta$ increases over the range $0\leq\Delta\lesssim 0.35$, the effects of increasing $\phi_{\rm J}$ (Fig. 1), decreasing $Z_{\rm J}$ (Fig. 2), and the fact that the sum in Eq. 8 is increasingly dominated by the energetically-costly contacts between large particles that dominate the jammed backbone Monti et al. (2022, 2023) may roughly cancel each other out. Alternatively, this trend could indicate that both $K_{\rm aff}$ and $K_{\rm na}$ increase with increasing $\Delta$, by roughly equal amounts, over this range. For $\Delta\gtrsim 0.35$, we observe that mean-squared nonaffine particle displacements (for any given $\Delta\phi$) increase rapidly with $\Delta$, suggesting that $K_{\rm na}$ also increases substantially with $\Delta$ foo . Alternatively, the negative $\partial K_{0}/\partial\Delta$ could arise primarily from the sharply-decreasing $Z_{\rm J}$.
@@
In systems prepared with SWAP and TDOF moves, the $\Delta$-dependence of $K_{0}$ is dramatically different. Instead of remaining nearly $\Delta$-independent over a wide range, $K_{0}$ increases rapidly over the range $0\leq\Delta\lesssim 0.05$ before plateauing over the range $0.05\lesssim\Delta\lesssim 0.10$. Next $\partial K_{0}/\partial\Delta$ increases (with a negative $\partial^{2}K_{0}/\partial\Delta^{2}$) over the range $0.10\lesssim\Delta\lesssim 0.20$. For $0.20\lesssim\Delta\lesssim 0.45$, $K_{0}$ increases linearly with $\Delta$ as indicated by the dashed line. Finally, for $\Delta\gtrsim 0.45$, $K_{0}$ increases above this linear trend. Clearly the trends in $K_{0}(\Delta)$ closely track the trends in $\phi_{\rm J}(\Delta)$ shown in Fig. 1, especially for $\Delta\gtrsim 0.20$. Other contributing factors include these packings’ much-larger $Z_{\rm J}(\Delta)$ (Fig. 2) and better-organized coordination shells (Fig. 5). These factors can act in concert to reduce nonaffine particle motion during sample compression, and therefore (presumably) to reduce $K_{\rm na}$ foo .
@@
The various trends discussed above combine to make the packings generated using SWAP and TDOF moves far more mechanically-stable. As shown in Fig. 6’s inset, the fractional increases in $K_{0}(\Delta)$ obtained by employing these moves increase rapidly from $\simeq 2\%$ to $\simeq 40\%$ over the range $0\leq\Delta\lesssim 0.05$, then increase more gradually (from $\simeq 40\%$ to $\simeq 70\%$) over the range $0.05\leq\Delta\lesssim 0.35$. For larger $\Delta$, the decreasing $K_{0}$ for systems prepared without these moves and increasing $\partial K_{0}/\partial\Delta$ for systems prepared with these moves produce even-more-dramatic fractional increases in $K_{0}$, reaching a maximum of 154% for $\Delta=0.50$. While these stability gains are somewhat smaller if $K_{0}$ is obtained by fitting the $P(\Delta\phi)$ data at lower strains – for example, the increase for $\Delta=0.50$ is only 80% if one employs the range $10^{-5}\leq\Delta\phi\leq 10^{-3}$ – all qualitative trends discussed above remain the same.

### v2

@@ III.1 Jamming densities and coordination numbers
In systems prepared without SWAP and TDOF moves, $Z_{\rm J}={\color[rgb]{1,0,0}5.91}$ for $\Delta=0$, and decreases monotonically with increasing $\Delta$, to a minimum of $Z_{\rm J}={\color[rgb]{1,0,0}3.13}$ for $\Delta=0.50$. Over the same range of $\Delta$, the rattler fraction increases monotonically from 3.8% to 63%. Both of these trends arise because (as discussed in Section II) our LS-like growth algorithm terminates and systems are considered jammed when $\tilde{\mathcal{G}}$ drops to zero within numerical precision, i.e. when the particles within the percolating, mechanically-rigid backbones whose formation normally coincides with jamming Moukarzel (1998) can no longer grow without producing overlaps. As $\Delta$ increases, these backbones are increasingly dominated by the large particles Monti et al. (2022, 2023), and thus the jammed states contain an increasing number of small rattlers. We emphasize, however, that these results are not artifacts of employing a LS-like particle-growth algorithm; the same trends are present in jammed Hookean-sphere packings prepared via compression at constant pressure Monti et al. (2022, 2023).
Adding SWAP moves to the packing-generation procedure yields far-higher $Z_{\rm J}(\Delta)$ for all $\Delta>0$. As $\Delta$ increases, $Z_{\rm J}$ first increases rapidly to a maximum of $6.81$ at $\Delta=0.04$, then decreases in a concave-up fashion over the range $0.04<\Delta\lesssim 0.25$, then decreases in a concave-down fashion, reaching its minimum value ($4.46$) at $\Delta=0.50$. For small polydispersities, these increases in $Z_{\rm J}$ are directly associated with dramatic reductions in the rattler fractions; we find $f_{\rm ratt}<0.01$ over the range $0.02\leq\Delta\leq 0.23$. On the other hand, SWAP moves rapidly lose their propensity to eliminate rattlers as $\Delta$ increases beyond $0.25$, and we find that employing them increases the rattler fractions for $\Delta>0.47$. Note that $f_{\rm ratt}$ usually increases with increasing $\phi_{\rm J}$ for fixed $\Delta$ Ozawa et al. (2017).
Systems prepared with SWAP and TDOF moves have far more contacts for both $\Delta=0$ and all $\Delta\gtrsim 0.25$. For $\Delta=0$, the TDOF moves increase $\Delta$ from $5.91$ to $6.09$. With increasing polydispersity, $Z_{\rm J}$ first passes through a local maximum ($Z_{\rm J}\simeq{\color[rgb]{1,0,0}6.82}$) at $\Delta={\color[rgb]{1,0,0}0.04}$, then decreases slowly before plateauing at 6.526.546.52-6.54 over the range $0.21\lesssim\Delta\lesssim 0.37$. Finally, for $\Delta>0.37$, $Z_{\rm J}$ increases steadily, to a maximum of $6.62$ at $\Delta=0.50$. As illustrated in the inset, adding the TDOF moves produces a further substantial reduction in $f_{\rm ratt}$ for both small and large polydispersity. The latter trend corresponds directly to the divergence at $\Delta\simeq 0.25$ of the $Z_{\rm J}(\Delta)$ curves for systems generated using both SWAP and TDOF moves and systems generated using only SWAP moves. We will argue below that this occurs because the TDOF moves allow smaller rattlers present in the system at the end of the first LS-like particle-growth stage (Section II) to continue growing until they also jam.
@@
The origin of $Z_{\rm J}(\Delta)$’s very-strong protocol-dependence is clarified by examining the probability distributions $P(Z_{i})$, i.e. the probabilities that a given particle has $Z_{i}$ contacts. Results are shown in Figure 3. Systems prepared without SWAP and TDOF moves have many “floaters” (particles with $Z_{i}=0$), and also substantial populations of rattlers with contacts (particles with $1\leq Z_{i}\leq 3$). Their $P(Z_{i})$ increase rapidly with increasing $\Delta$ for all $0\leq Z_{i}\leq 3$; similar results were reported in Refs. Oquendo-Patino and Estrada (2020, 2022). Minimally-mechanically-stable particles with $Z_{i}=4\equiv d+1$ (where $d$ is the spatial dimension non ) exhibit a qualitatively-different trend; $P(4)$ is nonomonotonic in $\Delta$, peaking at $\Delta\simeq{\color[rgb]{1,0,0}0.34}$.
Figure 3: Protocol dependence of coordination-number probability distributions $P(Z_{i})$ for $N=1000$ and $C=10^{-3}$. Solid and dashed curves show results obtained with and without SWAP and TDOF moves, while the vertical dashed lines indicate $Z_{i}=4$ and $Z_{i}=6$. Increasing/decreasing $C$ lengthens/shortens the tails of these distributions, but does not change the qualitative trends shown here.
@@
The shape of the $P(Z_{i})$ distributions also changes qualitatively as polydispersity increases. For all $\Delta\leq 0.13$, $P(6)>P(5)>P(4)$. Then, for $0.14\leq\Delta\leq 0.20$, $P(5)>P(6)>P(4)$. Next, for $0.21\leq\Delta\leq 0.26$, $P(5)>P(4)>P(6)$. Finally, for all $\Delta>0.26$, $P(4)>P(5)>P(6)$. This smooth crossover from dominance of locally-isostatic particles with $Z_{i}=6$ to minimally-mechanically-stable particles with $Z_{i}=4$ has not (to the best of our knowledge) been previously reported. The emergence of the peak at $Z_{i}=4$ coincides with the emergence of long high-$Z_{i}$ tails in the $P(Z_{i})$ that correspond to large spheres contacted by many smaller spheres; related issues were discussed in Refs. Monti et al. (2022, 2023); Srivastava et al. (2025).
@@
In systems prepared with SWAP and TDOF moves, the $P(Z_{i})$ are radically different. First, all floaters and rattlers with $Z_{i}\leq 2$ are eliminated. The remaining rattlers all have $Z_{i}=3$, and the $P(3)$ are at least three times lower than in systems prepared without these moves for all $\Delta>0$. The $P(4)$ are lower (higher) for all $\Delta<0.25$ ($\Delta\geq 0.25$). Finally, for all $Z_{i}>4$, the $P(Z_{i})$ are larger and decrease slower with increasing $Z_{i}$ than in systems prepared without these moves, particularly for $\Delta\gtrsim 0.2$.
All of the protocol-dependencies illustrated in Figs. 2-3 can be explained as follows. LS-like particle-growth algorithms, when performed without SWAP moves, terminate with large numbers of floaters/rattlers occupying the gaps between the particles forming the jammed backbone. For $\Delta\lesssim 0.25$, adding SWAP moves to the growth process eliminates most of these rattlers. For larger $\Delta$, SWAP moves are insufficient because many smaller particles remain free to rattle within cages formed by larger particles, but adding TDOF moves to the growth process allows these rattlers to continue growing until most of them also jam. We believe that the remaining rattlers indicate that the algorithm described in Sec. II is imperfect. Improving it might produce ideally-stable, rattler-free packings – analogous to those obtained in Ref. bol for a wider range of $\Delta$.
Relatively few studies of jamming have examined hyperstatic packings, i.e. systems with $Z_{\rm J}>Z_{\rm iso}\equiv 6$. Brito et al.  found that SWAP dynamics increase the number of contacts required to mechanically stabilize polydisperse packings, with $\delta Z=Z_{\rm J}-Z_{\rm iso}\sim\Delta^{1/2}$ Brito et al. (2018). Here we found that the combination of SWAP and TDOF moves produces substantailly-hyperstatic packings for all $\Delta>0$. The ($\phi_{\rm J},\Psi$)-related concepts discussed in Refs. Torquato et al. (2000); Donev et al. (2004, 2005); Brito et al. (2018); Anzivino et al. (2023) indicate that such hyperstatic packings must possess enhanced structural order, but they do not specify the form such order will take. Moreover, the connections between the hyperstaticity and mechanical stability of our higher-$\Delta$ packings are far less clear than is the case for their low-$\Delta$ counterparts. For example, we find that the fraction of particles with $Z_{i}>12$, which is impossible for monodisperse hard spheres, is nonzero for all $\Delta\geq 0.10$. This fraction exceeds 1% for all $\Delta>{\color[rgb]{1,0,0}0.26}$ ($\Delta>{\color[rgb]{1,0,0}0.19}$) in systems prepared without (with) SWAP and TDOF moves. The emergence of long tails in the $P(Z_{i})$ distributions for $\Delta\gtrsim 0.25$ indicates that caution must be exercised when using $Z_{\rm J}$ as a metric to assess high-$\Delta$ packings’ mechanical stability. In the following sections, we will explore these questions in detail.
@@
III.2 Other structural-order metrics
@@
One might expect that the changes in $Z_{\rm J}(\Delta)$ and $P(Z_{i})$ enabled by employing TDOF moves would necessarily coincide with substantial changes in the jammed states’ $P(\sigma)$ for $\Delta\gtrsim 0.25$. The latter would cast doubt on the utility of our results since TDOF moves are unphysical. However, as illustrated in Figure 4, the changes in $P(\sigma)$ associated with employing TDOF moves are negligible for all $0<\Delta\lesssim 0.45$. For $\Delta=0$, when the moves are not employed, $P(\sigma)=\delta(\sigma-1)$. When they are, $P(\sigma)$ is nonzero over a small range ${\color[rgb]{0,0,1}0.99}<\sigma<1.04$; this is only a very small broadening. For $\Delta\gtrsim 0.45$, $P(\sigma)$ decreases slightly for $\sigma\simeq\sigma_{\rm min}$ because the smallest particles are most able to grow as they fill in the jammed backbone’s gaps. We believe that this effect is responsible for the increase of $\phi_{\rm J}(\Delta)$ above its linear-fit value (Fig. 1).
@@
The TDOF moves employed in Refs. Hagh et al. (2022); bol produced much larger changes in $P(\sigma)$ because they were applied throughout the particle-growth process. In contrast, our approach (like that of Ref. Kim and Hilgenfeldt (2024)) minimizes changes in $P(\sigma)$ by not implementing TDOF moves until near the end of this process. Thus one can imagine that the effects of including TDOF moves highlighted in Fig. 2 could be at least partially reproduced in experiments, using a suitably designed packing-preparation protocol.
@@
Figure 5: Protocol dependence of pair correlation functions and structure factors for $N=1000$. Solid and dashed curves show results obtained with and without SWAP and TDOF moves, and the insets highlight the protocol-dependencies for $\Delta=0.10$ and $0.20$. All lengths have been scaled to yield $\langle\sigma\rangle=1$. Similar $\Delta$-dependence of $g(r)$ in packings prepared without SWAP was reported in Ref. Oquendo-Patino and Estrada (2020).
Figure 6: Typical jammed states for $N=1000$. The top and bottom rows respectively show packings generated without and with SWAP and TDOF moves. Columns from left to right show $\Delta=0.10,\ 0.20,\ 0.30,\ 0.40,\ \rm{and}\ 0.50$. Particle colors range from purple to red in order of increasing $\sigma$. Only spheres whose centers lie within the boxes are shown.
Previous studies of how preparation protocol affects jammed packings’ ($\phi_{\rm J},\Psi$) Donev et al. (2004); Ozawa et al. (2017) have shown that for fixed $P(\sigma)$ and $\Delta$, systems with larger $\phi_{\rm J}$ have higher rattler fractions $f_{\rm ratt}$, more bond-orientational and icosahedral-like order, reduced Voronoi-cell asphericity, and larger low-$k$ spectral densities (i.e. smaller long-wavelength packing-fraction fluctuations Zachary et al. (2011)). We have already shown that employing TDOF moves reverses the $\phi_{\rm J}$-dependence of $f_{\rm ratt}$, and we do not attempt to duplicate the other structural analyses here. Instead we ask a simpler question: does employing SWAP and TDOF moves during particle growth qualitatively change packings’ structural order?
@@
Examining the packings’ pair correlation functions $g(r)$ yields additional insights. As illustrated in Figure 5(a), employing SWAP and TDOF moves can produce packings with better-organized coordination shells. In particular, it yields a clearer separation between nearest and second-nearest neighbors, as indicated by the deeper first minima and higher second maxima of $g(r)$. These differences, however, only occur for $\Delta\lesssim{\color[rgb]{0,0.88,0}0.25}$. For $\Delta\gtrsim{\color[rgb]{0,0.88,0}0.25}$, the $g(r)$ are nearly protocol-independent. The increasingly-wide $P(\sigma)$ distributions (Fig. 4) make the first and second coordination shells indistinguishable, and the $g(r)$ are nearly ideal-gas-like for $r>\sigma_{\rm max}$.
Another useful metric for characterizing these packings’ order is the static structure factor $S(q)$. Fig. 5(b) shows that the main differences in $S(q)$ produced by employing SWAP and TDOF moves directly reflect the abovementioned differences in $g(r)$. More specifically, employing these moves increases the height of the first peaks in $S(q)$ and shifts them to larger $q$, but only for $\Delta\lesssim 0.2$. These trends directly correspond to the abovementioned deepening and shift to smaller $r$ of the first minima in $g(r)$. Another result apparent from this figure is that while fractionation increases with increasing $\Delta$ [as indicated by the increasing $\lim_{q\to 0}S(q)$], the degree to which this is so is protocol-independent. In other words, employing SWAP and TDOF moves does not seem to increase fractionation, as it did for hard-sphere liquids with $P(\sigma)\sim\sigma^{-3}$ and $\Delta\lesssim 0.10$ that were equilibrated at sufficiently large $\phi_{\rm eq}$ Ninarello et al. (2017).
Are these structural differences apparent to the naked eye? Figure 6 shows that the changes are, in fact, rather subtle. In particular, employing SWAP moves does not produce fractionated polycrystalline packings, as it did for low-$\alpha$ ellipses Hoy (2024). The only difference that is clear in these snapshots is that employing SWAP and TDOF moves allows small particles to more efficiently fill the gaps between the larger ones. Note, however, that we cannot rule out the presence of some local crystallization for small $\Delta$ near the peak in $Z_{\rm J}$ illustrated in Fig. 2.
III.3 Bulk moduli
@@
Finally, to demonstrate that the various structural differences discussed above strongly affect the packings’ mechanical properties, we examine their bulk moduli $K(\Delta)$. Multiple previous studies have examined how $K$ depends on the pair potential, particle size ratio (in bidisperse systems), and the overcompression $\Delta\phi=\phi-\phi_{\rm J}$ O’Hern et al. (2003); Goodrich et al. (2016); Petit et al. (2022); Mizuno et al. (2016). Here we focus on how $K$ depends on $\Delta$ and preparation protocol. For Hertzian interactions, $P\sim(\Delta\phi)^{3/2}$ in the small-$\Delta\phi$, elastic regime where systems deform affinely O’Hern et al. (2003). The resulting bulk moduli are singular at jamming, i.e. $K\sim(\Delta\phi)^{1/2}$, but a useful measure of their relative magnitudes can be obtained by writing
@@
Figure 7 shows the $K_{0}$ values obtained by fitting the $P(\Delta\phi)$ data to $P=P_{0}(\Delta\phi)^{3/2}$ over the power-law scaling regime, i.e. over the range $10^{-6}\leq\Delta\phi\leq 10^{-3}$, and then plugging the obtained $P_{0}$ values and results for $\phi_{\rm J}$ discussed above into Eq. 7.
Figure 7: Protocol dependence of bulk moduli for $N=1000$. Panel (a) shows results for $P(\Delta\phi)$, while panel (b) shows our estimates for $K_{0}$. In panel (a), solid and dashed curves show results obtained with and without SWAP and TDOF moves, and the dot-dashed lines indicate $P\sim(\Delta\phi)^{3/2}$ scaling. In panel (b), rhe dashed line shows a linear fit to $K_{0}$ over the range $0.25\leq\Delta\leq 0.45$, and the inset shows the fractional increase in $K_{0}$ obtained by employing SWAP and TDOF moves. All lengths have been scaled to yield $\langle\sigma\rangle=1$, and $K_{0}$ values are given in urits of $\varepsilon/\langle\sigma\rangle^{3}$.
@@
is the classical Born term arising from affine compression and $K_{\rm na}$ captures the reduction of $K$ by nonaffine particle displacements Maloney and Lemaître (2004); Zaccone and Scossa-Romano (2011); Mizuno et al. (2016); Milkus and Zaccone (2016). The $1/V$ prefactor and the purely-repulsive interactions [i.e., $U(r_{ij})=0$ for $r_{ij}>\sigma_{ij}$] respectively imply $K\sim\phi_{\rm J}$ and $K\sim Z_{\rm J}$ for monodisperse systems. Polydisperse systems are more complicated, as the $Z_{\rm J}$-dependence of their bulk moduli depends sensitively on which particles are in contact. Our results suggest that as $\Delta$ increases over the range $0\leq\Delta\lesssim 0.35$, the effects of increasing $\phi_{\rm J}$ (Fig. 1), decreasing $Z_{\rm J}$ (Fig. 2), and the fact that the sum in Eq. 8 is increasingly dominated by the energetically-costly contacts between large particles that dominate the jammed backbone Monti et al. (2022, 2023) may roughly cancel each other out.
@@
In systems prepared with SWAP and TDOF moves, the $\Delta$-dependence of $K_{0}$ is dramatically different. Instead of remaining nearly $\Delta$-independent over a wide range, $K_{0}$ increases rapidly over the range $0\leq\Delta\lesssim{\color[rgb]{0,0,1}0.03}$ before plateauing over the range $0.03\lesssim\Delta\lesssim 0.12$. Next $\partial K_{0}/\partial\Delta$ increases (with a negative $\partial^{2}K_{0}/\partial\Delta^{2}$) over the range $0.12\lesssim\Delta\lesssim 0.25$. For ${\color[rgb]{0,0,1}0.25}\lesssim\Delta\lesssim 0.45$, $K_{0}$ increases linearly with $\Delta$ as indicated by the dashed line. Finally, for $\Delta\gtrsim 0.45$, $K_{0}$ increases above this linear trend. Clearly these trends in $K_{0}(\Delta)$ closely track the trends in $\phi_{\rm J}(\Delta)$ shown in Fig. 1, especially for $\Delta\gtrsim{\color[rgb]{0,0,1}0.25}$.
@@
We believe that the higher $\phi_{\rm J}(\Delta)$ (Fig. 1), larger $Z_{\rm J}(\Delta)$ (Fig. 2) and better-organized coordination shells (Fig. 5) combine to make the packings generated using SWAP and TDOF moves far more mechanically-stable. As shown in Fig. 7’s inset, the fractional increases in $K_{0}(\Delta)$ obtained by employing these moves increase rapidly from $\simeq 3\%$ to $\simeq 25\%$ over the range $0\leq\Delta\lesssim{\color[rgb]{0,0,1}0.03}$, then increase more gradually (from $\simeq 25\%$ to $\simeq 45\%$) over the range ${\color[rgb]{0,0,1}0.03}\leq\Delta\lesssim 0.35$. For larger $\Delta$, the decreasing $K_{0}$ for systems prepared without these moves and increasing $\partial K_{0}/\partial\Delta$ for systems prepared with these moves produce even-more-dramatic fractional increases in $K_{0}$, reaching a maximum of 80% for $\Delta=0.50$.
Here, since we are focusing mainly on structure rather than mechanics, we presented only a brief analysis of systems’ bulk moduli, to illustrate the basic result that polydisperse 3D sphere packings generated using SWAP and TDOF moves are far more mechanically-stable than packings generated without these moves, and that the magnitude of these stability increases grows rapidly with increasing $\Delta$. We have chosen not to speculate about the role nonaffine deformation plays in the above results because rigorously testing any hypotheses about $K_{\rm na}$ would require detailed analyses Maloney and Lemaître (2004); Zaccone and Scossa-Romano (2011); Mizuno et al. (2016); Milkus and Zaccone (2016) that are beyond our present scope. We plan to explore this issue further in a followup study that wlll characterize these packings’ linear and nonlinear mechanics, under both compression and shear, in far more detail.